Sol Summary – April 1st

Sol: 1
Summary Title: Getting used to Mars
Author’s name: Alba Sánchez Montalvo
Mission Status: initiated
Sol Activity Summary:
First day on Mars! This morning, the crew ATLAS had breakfast together after taking their probiotic/placebo treatment and prepared for their morning EVA which was mainly about drones. Half of the crew performed the baseline register for Romain’s experiment and Louis started to evaluate the surroundings for his experiment. During this time, part of the team also started the food summary and Maxime built his weather station mount in the RAM. After lunch, Alba, Arnaud and Imane performed blood cell count in a few drops of blood and collected saliva from the crew members for their biomedical research. In the afternoon, the other half of the team went on an EVA to do the baseline of Romain’s experiment. Once they were back to the station, Alba collected some blood from the crew members to extract serum and they also performed a salivary test for their studies. The crew had a good time together at the science dome during the analysis and right after, they came back to the main hab and they filled some forms for Loriane and Arnaud’s experiment on the psychological part of the simulation. During the day, everyone collected their feces and started collecting urines for the analysis of the biomedical team. To finish the day, the crew worked together on the daily reports.
Look Ahead Plan: the crew will do an EVA tomorrow. The purpose of the EVA involves two main experiments conducted by four crew members. Maxime and Loriane will work on setting up a weather station near the hab using the Curiosity rover, while Louis and Hippolyte will utilize the Spirit rover for Louis’ experiment, which involves conducting automated flights with a drone using PIX4Dcapture software. Both experiments focus on exploring the area around the hab, particularly the North Ridge, to gather data on weather conditions, terrain, and the feasibility of automated flight missions for future exploration.
Anomalies in work: Romain’s drone lost its connection during the EVA in the afternoon. Louis forgot some material for the complete use of his drone.
Weather: Cloudy and sunny in the morning, a little windy and warmer in the afternoon.
Crew Physical Status: nothing special, everyone feels good and active.
EVA: Despite Romain’s issue with the drone, we achieved our planned goals.
Reports to be filed: Sol summary, Journalist’s report, Greenhab report, operations report, EVA report, EVA request, astronomy report.
Support Requested: we asked for a table to place it in the science dome.

Copyright © The Mars Society. All rights reserved. | Main Site